8. find the vertex, axis of symmetry and graph. f(x) = 2(x - 2)2 - 3.
Answer #1

Sot - Given that f(x)=2(2-2) - 3 compare egs 0 with f(x) = 2(x-btk where vester (hk) & The vertex of f(x) .(bk) = ( 2-3) & ve= -50 (4,5) (0,5) -5 0 5 (2, -3) + « Х f(x) = 2(x - 2)2 – 3 2 X (2,-3) Label: 3 (0,5) Х Label: 4 (4,5) Х Label: 5 (5,15) Х La